# Optimal Multi-Distribution Learning

### v2

@@ Abstract
Multi-distribution learning (MDL), which seeks to learn a shared model that minimizes the worst-case risk across $k$ distinct data distributions, has emerged as a unified framework in response to the evolving demand for robustness, fairness, multi-group collaboration, etc. Achieving data-efficient MDL necessitates adaptive sampling, also called on-demand sampling, throughout the learning process. However, there exist substantial gaps between the state-of-the-art upper and lower bounds on the optimal sample complexity. Focusing on a hypothesis class of Vapnik–Chervonenkis (VC) dimension $d$, we propose a novel algorithm that yields an $\varepsilon$-optimal randomized hypothesis with a sample complexity on the order of $\frac{d+k}{\varepsilon^{2}}$ (modulo some logarithmic factor), matching the best-known lower bound. Our algorithmic ideas and theory have been further extended to accommodate Rademacher classes. The proposed algorithms are oracle-efficient, which access the hypothesis class solely through an empirical risk minimization oracle. Additionally, we establish the necessity of randomization, unveiling a large sample size barrier when only deterministic hypotheses are permitted. These findings successfully resolve three open problems presented in COLT 2023 (i.e., Awasthi et al., (2023, Problems 1, 3 and 4)).
@@ 1 Introduction
Driven by the growing need of robustness, fairness and multi-group collaboration in machine learning practice, the multi-distribution learning (MDL) framework has emerged as a unified solution in response to these evolving demands (Blum et al.,, 2017; Haghtalab et al.,, 2022; Mohri et al.,, 2019; Awasthi et al.,, 2023). Setting the stage, imagine that we are interested in a collection of $k$ unknown data distributions $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{k}$ supported on $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ (resp. $\mathcal{Y}$) stands for the instance (resp. label) space. Given a hypothesis class $\mathcal{H}$ and a prescribed loss function [^1] For example, for each hypothesis $h\in\mathcal{H}$ and each datapoint $(x,y)\in\mathcal{X}\times\mathcal{Y}$, we employ $\ell(h,(x,y))$ to measure the risk of using hypothesis $h$ to predict $y$ based on $x$. :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$, we are asked to identify a (possibly randomized) hypothesis $\widehat{h}$ achieving near-optimal worst-case loss across these data distributions, namely, [^2] Here, the expectation on the left-hand side of (1) is taken over the randomness of both the datapoints $(x,y)$ and the (randomized) hypothesis $\widehat{h}$.
@@
with $\varepsilon\in(0,1]$ a target accuracy level. In light of the unknown nature of these data distributions, the learning process involves data collection, allowing the learner to sample from $\{\mathcal{D}_{i}\}_{i=1}^{k}$. The performance of a learning algorithm is then gauged by its sample complexity — the number of samples required to fulfil (1). Our objective is to design a learning paradigm that achieves optimal sample complexity.
@@
The MDL framework described above, which can viewed as an extension of agnostic learning (Valiant,, 1984; Blumer et al.,, 1989) tailored to multiple data distributions, has found a wealth of applications across multiple domains. Here, we highlight a few representative examples, and refer the interested reader to Haghtalab et al., (2022) for more in-depth discussions.
@@
  • Collaborative and agnostic federated learning. In the realm of collaborative and agnostic federated learning (Blum et al.,, 2017; Nguyen and Zakynthinou,, 2018; Chen et al.,, 2018; Mohri et al.,, 2019; Blum et al., 2021a, ; Du et al.,, 2021; Deng et al.,, 2020; Blum et al., 2021b, ), a group of $k$ agents, each having access to distinct data sources as characterized by different data distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$, aim to learn a shared prediction model that ideally would achieve low risk for each of their respective data sources. A sample-efficient MDL paradigm would help unleash the potential of collaboration and information sharing in jointly learning a complicated task.
@@
  • Min-max fairness in learning. The MDL framework is well-suited to scenarios requiring fairness across multiple groups (Dwork et al.,, 2021; Rothblum and Yona,, 2021; Du et al.,, 2021). For instance, in situations where multiple subpopulations with distinct data distributions exist, a prevailing objective is to ensure that the learned model does not adversely impact any of these subpopulations. Such a criterion, known as “min-max fairness” in the literature (Mohri et al.,, 2019; Abernethy et al.,, 2022), plays a pivotal role in mitigating the worst disadvantage experienced by any particular subpopulation.
@@
  • Distributionally robust optimization/learning. Another context where MDL naturally finds applications is group distributionally robust optimization and learning (DRO/DRL). Group DRO and DRL aim to develop algorithms that offer robust performance guarantees across a finite number of possible distributional models (Sagawa et al.,, 2019, 2020; Hashimoto et al.,, 2018; Hu et al.,, 2018; Xiong et al.,, 2023; Zhang et al.,, 2020; Wang et al.,, 2023; Deng et al.,, 2020), and have garnered substantial attention recently due to the pervasive need for robustness in modern decision-making applications (Carmon and Hausler,, 2022; Asi et al.,, 2021; Haghtalab et al.,, 2022; Kar et al.,, 2019). When applying MDL to the context of group DRO/DRL, the resultant sample complexity reflects the price that needs to be paid for learning a robust solution.
@@
The MDL framework is also closely related to other topics like multi-source domain adaptation, maximum aggregation, to name just a few (Mansour et al.,, 2008; Zhao et al.,, 2020; Bühlmann and Meinshausen,, 2015; Guo,, 2023).
@@
In stark contrast to single-distribution learning, achieving data-efficient MDL necessitates adaptive sampling throughout the learning process, also known as on-demand sampling (Haghtalab et al.,, 2022). More specifically, pre-determining a sample-size budget for each distribution beforehand and sampling non-adaptively could result in a loss of sample efficiency, as we lack knowledge about the complexity of learning each distribution before the learning process begins. The question then boils down to how to optimally adapt the online sampling strategy to effectively tackle diverse data distributions.
@@
The sample complexity of MDL has been explored in a strand of recent works under various settings. Consider first the case where the hypothesis class $\mathcal{H}$ comprises a finite number of hypotheses. If we sample non-adaptively and draw the same number of samples from each individual distribution $\mathcal{D}_{i}$, then this results in a total sample size exceeding the order of $\frac{k\log(|\mathcal{H}|)}{\varepsilon^{2}}$ (given that learning each distribution requires a sample size at least on the order of $\frac{\log(|\mathcal{H}|)}{\varepsilon^{2}}$). Fortunately, this sample size budget can be significantly reduced with the aid of adaptive sampling. In particular, the state-of-the-art approach, proposed by Haghtalab et al., (2022), accomplishes the objective (1) with probability $1-\delta$ using $O\big{(}\frac{\log(|\mathcal{H}|)+k\log(k/\delta)}{\varepsilon^{2}}\big{)}$ samples. In comparison to agnostic learning on a single distribution, it only incurs an extra additive cost of $k\log(k/\delta)/\varepsilon^{2}$ as opposed to a multiplicative factor in $k$, thus underscoring the importance of adaptive sampling.
A more challenging scenario arises when $\mathcal{H}$ has a finite Vapnik–Chervonenkis (VC) dimension $d$. The sample complexity for VC classes has only been settled for the reliazable case (Blum et al.,, 2017; Chen et al.,, 2018; Nguyen and Zakynthinou,, 2018), a special scenario where the loss function takes the form of $\ell\big{(}h,(x,y)\big{)}=\mathds{1}\{h(x)\neq y\}$ and it is feasible to achieve zero mean loss. For the general non-realizable case, the best-known lower bound for such VC classes is (Haghtalab et al.,, 2022) [^3] Here and throughout, $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ hide all logarithimic factors in $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
@@
which serves as a theoretical benchmark. By first collecting $\widetilde{O}\left(\frac{dk}{\varepsilon}\right)$ samples to help construct a cover of $\mathcal{H}$ with reasonable resolution, Haghtalab et al., (2022) established a sample complexity upper bound of
@@
Nevertheless, the term $dk/\varepsilon$ in (3a) fails to match the lower bound (2); put another way, this term might result in a potentially large burn-in cost, as the optimality of this approach is only guaranteed (up to log factors) when the total sample size already exceeds an enormous threshold on the order of $\frac{d^{2}k^{2}}{d+k}$. In an effort to alleviate this $dk/\varepsilon$ factor, Awasthi et al., (2023) put forward an alternative algorithm — which utilizes an oracle to learn on a single distribution and obliviates the need for computing an epsilon-net of $\mathcal{H}$ — yielding a sample complexity of
@@
Notably, this question has been posed as an open problem during the Annual Conference on Learning Theory (COLT) 2023; see Awasthi et al., (2023, Problem 1).
@@
Haghtalab et al., (2022) $\frac{\log(|\mathcal{H}|)+k}{\varepsilon^{2}}$
Haghtalab et al., (2022) $\frac{d+k}{\varepsilon^{2}}+\frac{dk}{\varepsilon}$
Awasthi et al., (2023) $\frac{d}{\varepsilon^{4}}+\frac{k}{\varepsilon^{2}}$
Peng, (2023) $\frac{d+k}{\varepsilon^{2}}\left(\frac{k}{\epsilon}\right)^{o(1)}$
@@
lower bound: Haghtalab et al., (2022) $\frac{d+k}{\varepsilon^{2}}$
Table 1: Sample complexity bounds of MDL with $k$ data distributions and a hypothesis class of VC dimension $d$. Here, we only report the polynomial depedency and hide all logarithmic dependency on $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
@@
The present paper delivers some encouraging news: we come up with a new MDL algorithm that successfully resolves the aforementioned open problem in the affirmative. Specifically, focusing on a hypothesis class with VC dimension $d$ and a collection of $k$ data distributions, our main findings can be summarized as follows.
@@
The polylog factor in (4) will be specified momentarily. In a nutshell, we develop the first algorithm that provably achieves a sample complexity matching the lower bound (2) modulo logarithmic factors. Following the game dynamics template adopted in previous methods — namely, viewing MDL as a game between the learner (who selects the best hypothesis) and the adversary (who chooses the most challenging mixture of distributions) — our algorithm is built upon a novel and meticulously designed sampling scheme that deviates significantly from previous methods. Further, we extend our algorithm and theory to accommodate Rademacher classes, establishing a similar sample complexity bound when the weighted Rademacher complexity of the hypothesis class on $n$ points is upper bounded by $O\big{(}\sqrt{\frac{d\log(n)}{n}}\big{)}$.
Additionally, we solve two other open problems posed by Awasthi et al., (2023):
  • Oracle-efficient solutions. An algorithm is said to be oracle-efficient if it only accesses $\mathcal{H}$ through an empirical risk minimization (ERM) oracle (Dudík et al.,, 2020). Awasthi et al., (2023, Problem 4) then asks what the sample complexity of MDL is when confined to oracle-efficient paradigms. Encouragingly, our algorithm (i.e., Algorithm 1) adheres to the oracle-efficient criterion, thus uncovering that the sample complexity of MDL remains unchanged when restricted to oracle-efficient algorithms.
  • Necessity of randomization. Both our algorithm and the most sample-efficient methods preceding our work produce randomized hypotheses. As discussed around Awasthi et al., (2023, Problem 3), a natural question concerns characterization of the sample complexity when restricting the final output to deterministic hypotheses from $\mathcal{H}$. Our result (see Theorem 18) delivers a negative message: under mild conditions, for any MDL algorithm, there exists a hard problem instance such that it requires at least $\Omega(dk/\varepsilon^{2})$ samples to find a deterministic hypothesis $h\in\mathcal{H}$ that attains $\varepsilon$-accuracy. This constitutes a sample size barrier that lies far above what is achievable under randomized hypotheses.
@@
While finalizing our paper, we became aware of a concurrent work Peng, (2023), which also studied the MDL problem and significantly improved upon the prior results. More specifically, Peng, (2023) established a sample complexity of $O\left(\frac{(d+k)\log(d/\delta)}{\varepsilon^{2}}\left(\frac{k}{\varepsilon}\right)^{o(1)}\right)$, which is optimal up to some sub-polynomial factor in $k/\varepsilon$; in comparison, our sample complexity is optimal only up to some logarithmic factor. Additionally, it is worth noting that the algorithm therein relies upon a certain recursive structure to eliminate the non-optimal hypothesis, thus incurring exponential computational cost even when an ERM oracle is available.
@@
Throughout this paper, we denote $[N]\coloneqq\{1,\ldots,N\}$ for any positive integer $N$. Let $\mathsf{conv}(\mathcal{A})$ represent the convex hull of a set $\mathcal{A}$, and denote by $\Delta(n)$ the $n$-dimensional simplex for any positive integer $n$. For two vectors $v=[v_{i}]_{1\leq i\leq n}$ and $v^{\prime}=[v^{\prime}_{i}]_{1\leq i\leq n}$ with the same dimension, we overload the notation by using $\max\{v,v^{\prime}\}=\big{[}\max\{v_{i},v^{\prime}_{i}\}\big{]}_{1\leq i\leq n}$ to denote the coordinate-wise maximum of $v$ and $v^{\prime}$. Also we say $v\leq v^{\prime}$ iff $v_{i}\leq v^{\prime}_{i}$ for all $i\in[n]$. For any random variable $X$, we use $\mathbb{V}[X]$ to denote its variance, i.e., $\mathbb{V}[X]=\mathbb{E}\left[(X-\mathbb{E}[X])^{2}\right]$. For any two distributions $P$ and $Q$ supported on $\mathcal{X}$, the Kullback-Leibler (KL) divergence from $Q$ to $P$ is defined and denoted by

### v3

@@ Abstract
Multi-distribution learning (MDL), which seeks to learn a shared model that minimizes the worst-case risk across $k$ distinct data distributions, has emerged as a unified framework in response to the evolving demand for robustness, fairness, multi-group collaboration, etc. Achieving data-efficient MDL necessitates adaptive sampling, also called on-demand sampling, throughout the learning process. However, there exist substantial gaps between the state-of-the-art upper and lower bounds on the optimal sample complexity. Focusing on a hypothesis class of Vapnik–Chervonenkis (VC) dimension $d$, we propose a novel algorithm that yields an $\varepsilon$-optimal randomized hypothesis with a sample complexity on the order of $\frac{d+k}{\varepsilon^{2}}$ (modulo some logarithmic factor), matching the best-known lower bound. Our algorithmic ideas and theory are further extended to accommodate Rademacher classes. The proposed algorithms are oracle-efficient, which access the hypothesis class solely through an empirical risk minimization oracle. Additionally, we establish the necessity of randomization, revealing a large sample size barrier when only deterministic hypotheses are permitted. These findings resolve three open problems presented in COLT 2023 (i.e., awasthi2023sample).
@@ 1 Introduction
Driven by the growing need of robustness, fairness and multi-group collaboration in machine learning practice, the multi-distribution learning (MDL) framework has emerged as a unified solution in response to these evolving demands (blum2017collaborative; haghtalab2022demand; mohri2019agnostic; awasthi2023sample). Setting the stage, imagine that we are interested in a collection of $k$ unknown data distributions $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{k}$ supported on $\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ (resp. $\mathcal{Y}$) stands for the instance (resp. label) space. Given a hypothesis class $\mathcal{H}$ and a prescribed loss function [^1] For example, for each hypothesis $h\in\mathcal{H}$ and each datapoint $(x,y)\in\mathcal{X}\times\mathcal{Y}$, we employ $\ell(h,(x,y))$ to measure the risk of using hypothesis $h$ to predict $y$ based on $x$. :×$\ell:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to[-1,1]$, we are asked to identify a (possibly randomized) hypothesis $\widehat{h}$ achieving near-optimal worst-case loss across these data distributions, namely, [^2] Here, the expectation on the left-hand side of (1) is taken over the randomness of both the datapoints $(x,y)$ and the (randomized) hypothesis $\widehat{h}$.
@@
with $\varepsilon\in(0,1]$ a target accuracy level. In light of the unknown nature of these data distributions, the learning process is often coupled with data collection, allowing the learner to sample from $\{\mathcal{D}_{i}\}_{i=1}^{k}$. The performance of a learning algorithm is then gauged by its sample complexity — the number of samples required to fulfil (1). Our objective is to design a learning paradigm that achieves the optimal sample complexity.
@@
The MDL framework described above, which can viewed as an extension of agnostic learning (valiant1984theory; blumer1989learnability) tailored to multiple data distributions, has found a wealth of applications across multiple domains. Here, we highlight a few representative examples, and refer the interested reader to haghtalab2022demand and the references therein for more extensive discussions.
@@
  • Collaborative and agnostic federated learning. In the realm of collaborative and agnostic federated learning (blum2017collaborative; nguyen2018improved; chen2018tight; mohri2019agnostic; blum2021one; du2021fairness; deng2020distributionally; blum2021communication), a group of $k$ agents, each having access to distinct data sources as characterized by different data distributions $\{\mathcal{D}_{i}\}_{i=1}^{k}$, aim to learn a shared prediction model that ideally would achieve low risk for each of their respective data sources. A sample-efficient MDL paradigm would help unleash the potential of collaboration and information sharing in jointly learning a complicated task.
@@
  • Min-max fairness in learning. The MDL framework is well-suited to scenarios requiring fairness across multiple groups (dwork2021outcome; rothblum2021multi; du2021fairness). For instance, in situations where multiple subpopulations with distinct data distributions exist, a prevailing objective is to ensure that the learned model does not adversely impact any of these subpopulations. One criterion designed to meet this objective, known as “min-max fairness” in the literature (mohri2019agnostic; abernethy2020active), plays a pivotal role in mitigating the worst disadvantage experienced by any particular subpopulation.
@@
  • Distributionally robust optimization/learning. Another context where MDL naturally finds applications is group distributionally robust optimization and learning (DRO/DRL). Group DRO and DRL aim to develop algorithms that offer robust performance guarantees across a finite number of possible distributional models (sagawa2019distributionally; sagawa2020investigation; hashimoto2018fairness; hu2018does; xiong2023distributionally; zhang2020coping; wang2023distributionally; deng2020distributionally), and have garnered substantial attention recently due to the pervasive need for robustness in modern decision-making (carmon2022distributionally; asi2021stochastic; haghtalab2022demand; kar2019meta). When applying MDL to the context of group DRO/DRL, the resultant sample complexity reflects the price that needs to be paid for learning a robust solution.
@@
The MDL framework is also closely related to other topics like multi-source domain adaptation, maximum aggregation, to name just a few (mansour2008domain; zhao2020multi; buhlmann2015magging; guo2023statistical).
@@
In contrast to single-distribution learning, achieving data-efficient MDL necessitates adaptive sampling throughout the learning process, also known as on-demand sampling (haghtalab2022demand). More specifically, pre-determining a sample-size budget for each distribution beforehand and sampling non-adaptively could result in loss of sample efficiency, as we lack knowledge about the complexity of learning each distribution before the learning process begins. The question then comes down to how to optimally adapt the online sampling strategy to effectively tackle diverse data distributions.
@@
The sample complexity of MDL has been explored in a strand of recent works under various settings. Consider first the case where the hypothesis class $\mathcal{H}$ comprises a finite number of hypotheses. If we sample non-adaptively and draw the same number of samples from each individual distribution $\mathcal{D}_{i}$, then this results in a total sample size exceeding the order of $\frac{k\log(|\mathcal{H}|)}{\varepsilon^{2}}$ (given that learning each distribution requires a sample size at least on the order of $\frac{\log(|\mathcal{H}|)}{\varepsilon^{2}}$). Fortunately, this sample size budget can be significantly reduced with the aid of adaptive sampling. In particular, the state-of-the-art approach, proposed by haghtalab2022demand, accomplishes the objective (1) with probability at least $1-\delta$ using $O\big{(}\frac{\log(|\mathcal{H}|)+k\log(k/\delta)}{\varepsilon^{2}}\big{)}$ samples. In comparison to agnostic learning on a single distribution, it only incurs an extra additive cost of $k\log(k/\delta)/\varepsilon^{2}$ as opposed to a multiplicative factor in $k$, thus underscoring the importance of adaptive sampling.
A more challenging scenario arises when $\mathcal{H}$ has a finite Vapnik–Chervonenkis (VC) dimension $d$. The sample complexity for VC classes has only been settled for the realizable case (blum2017collaborative; chen2018tight; nguyen2018improved), a special scenario where the loss function takes the form of $\ell\big{(}h,(x,y)\big{)}=\mathds{1}\{h(x)\neq y\}$ and it is feasible to achieve zero mean loss. For the general non-realizable case, the best-known lower bound for such VC classes is (haghtalab2022demand) [^3] Let $\mathcal{X}=\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$. Here and throughout, the notation $f(\mathcal{X})=O\big{(}g(\mathcal{X})\big{)}$ or $f(\mathcal{X})\lesssim g(\mathcal{X})$ (resp. $f(\mathcal{X})=\Omega\big{(}g(\mathcal{X})\big{)}$ or $f(\mathcal{X})\gtrsim g(\mathcal{X})$) mean that there exists some universal constant $C_{1}>0$ (resp. $C_{2}>0$) such that $f(\mathcal{X})\leq C_{1}g(\mathcal{X})$ (resp. $f(\mathcal{X})\geq C_{2}g(\mathcal{X})$); the notation $f(\mathcal{X})=\Theta\big{(}g(\mathcal{X})\big{)}$ or $f(\mathcal{X})\asymp g(\mathcal{X})$ mean $f(\mathcal{X})=O\big{(}g(\mathcal{X})\big{)}$ and $f(\mathcal{X})=\Omega\big{(}g(\mathcal{X})\big{)}$ hold simultaneously. The notation $\widetilde{O}(\cdot)$, $\widetilde{\Theta}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ are defined analogously except that all log factors in $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$ are hidden.
@@
which serves as a theoretical benchmark. By first collecting $\widetilde{O}\left(\frac{dk}{\varepsilon}\right)$ samples to help construct a cover of $\mathcal{H}$ with reasonable resolution, haghtalab2022demand established a sample complexity upper bound of
@@
Nevertheless, the term $dk/\varepsilon$ in (3a) fails to match the lower bound (2); put another way, this term might result in a potentially large burn-in cost, as the optimality of this approach is only guaranteed (up to log factors) when the total sample size already exceeds a (potentially large) threshold on the order of $\frac{d^{2}k^{2}}{d+k}$. In an effort to alleviate this $dk/\varepsilon$ factor, awasthi2023sample put forward an alternative algorithm — which utilizes an oracle to learn on a single distribution and obliviates the need for computing an epsilon-net of $\mathcal{H}$ — yielding a sample complexity of
@@
Notably, this question has been posed as an open problem during the Annual Conference on Learning Theory (COLT) 2023; see awasthi2023sample.
@@
haghtalab2022demand $\frac{\log(|\mathcal{H}|)+k}{\varepsilon^{2}}$
haghtalab2022demand $\frac{d+k}{\varepsilon^{2}}+\frac{dk}{\varepsilon}$
awasthi2023sample $\frac{d}{\varepsilon^{4}}+\frac{k}{\varepsilon^{2}}$
peng2023sample $\frac{d+k}{\varepsilon^{2}}\left(\frac{k}{\varepsilon}\right)^{o(1)}$
@@
lower bound: haghtalab2022demand $\frac{d+k}{\varepsilon^{2}}$
Table 1: Sample complexity bounds of MDL with $k$ data distributions and a hypothesis class of VC dimension $d$. Here, we only report the polynomial dependency and hide all logarithmic dependency on $\big{(}k,d,\frac{1}{\varepsilon},\frac{1}{\delta}\big{)}$.
@@
The present paper delivers some encouraging news: we come up with a new MDL algorithm that successfully resolves the aforementioned open problem in the affirmative. Specifically, focusing on a hypothesis class with VC dimension $d$ and a collection of $k$ data distributions, our main findings can be summarized in the following theorem. [^4] Following the definition of the VC dimension, the hypotheses in $\mathcal{H}$ are assumed to be binary-valued for VC classes.
@@
The polylog factor in (4) will be specified momentarily. In a nutshell, we develop the first algorithm that provably achieves a sample complexity matching the lower bound (2) modulo logarithmic factors. Following the game dynamics template adopted in previous methods — namely, viewing MDL as a game between the learner (who selects the best hypothesis) and the adversary (who chooses the most challenging mixture of distributions) — our algorithm is built upon a novel and meticulously designed sampling scheme that deviates significantly from previous methods. Further, we extend our algorithm and theory to accommodate Rademacher classes, establishing a similar sample complexity bound; see Section 7 for details.
Additionally, we solve two other open problems posed by awasthi2023sample:
  • Oracle-efficient solutions. An algorithm is said to be oracle-efficient if it only accesses $\mathcal{H}$ through an empirical risk minimization (ERM) oracle (dudik2020oracle). awasthi2023sample then asked what the sample complexity of MDL is when confined to oracle-efficient paradigms. Encouragingly, our algorithm (i.e., Algorithm 1) adheres to the oracle-efficient criterion, thus uncovering that the sample complexity of MDL remains unchanged when restricted to oracle-efficient algorithms.
  • Necessity of randomization. Both our algorithm and the most sample-efficient methods preceding our work produce randomized hypotheses. As discussed around awasthi2023sample, a natural question concerns characterization of the sample complexity when restricting the final output to deterministic hypotheses from $\mathcal{H}$. Our result (see Theorem 2) delivers a negative message: under mild conditions, for any MDL algorithm, there exists a hard problem instance such that it requires at least $\Omega(dk/\varepsilon^{2})$ samples to find a deterministic hypothesis $h\in\mathcal{H}$ that attains $\varepsilon$-accuracy. This constitutes an enormous sample complexity gap between what is achievable under randomized hypotheses and what is achievable using deterministic hypotheses.
@@
We shall mention that a concurrent work peng2023sample, posted around the same time as our work, also studied the MDL problem and significantly improved upon the prior results. More specifically, peng2023sample established a sample complexity of $O\left(\frac{(d+k)\log(d/\delta)}{\varepsilon^{2}}\left(\frac{k}{\varepsilon}\right)^{o(1)}\right)$, which is optimal up to some sub-polynomial factor in $k/\varepsilon$; in comparison, our sample complexity is optimal up to polylogarithmic factor. Additionally, it is worth noting that the algorithm therein relies upon a certain recursive structure to eliminate the non-optimal hypothesis, thus incurring exponential computational cost even when an ERM oracle is available.
@@
Throughout this paper, we denote $[N]\coloneqq\{1,\ldots,N\}$ for any positive integer $N$. Let $\mathsf{conv}(\mathcal{A})$ represent the convex hull of a set $\mathcal{A}$, and denote by $\Delta(n)$ the $n$-dimensional simplex for any positive integer $n$. For two vectors $v=[v_{i}]_{1\leq i\leq n}$ and $v^{\prime}=[v^{\prime}_{i}]_{1\leq i\leq n}$ with the same dimension, we overload the notation by using $\max\{v,v^{\prime}\}=\big{[}\max\{v_{i},v^{\prime}_{i}\}\big{]}_{1\leq i\leq n}$ to denote the coordinate-wise maximum of $v$ and $v^{\prime}$. Also we say $v\leq v^{\prime}$ iff $v_{i}\leq v^{\prime}_{i}$ for all $i\in[n]$. For any random variable $X$, we use $\mathsf{Var}[X]$ to denote its variance, i.e., $\mathsf{Var}[X]=\mathbb{E}\left[(X-\mathbb{E}[X])^{2}\right]$. Let $\mathrm{e}^{\mathsf{basis}}_{1},\dots,\mathrm{e}^{\mathsf{basis}}_{k}$ represent the standard basis vectors in $\mathbb{R}^{k}$. For any two distributions $P$ and $Q$ supported on $\mathcal{X}$, the Kullback-Leibler (KL) divergence from $Q$ to $P$ is defined and denoted by